# Integration of a high-fidelity model of quantum sensors with a map-matching filter for quantum-enhanced navigation

### v1

@@ Appendix B Estimation of systematic effects and mitigation requirements
We consider here a simplified single-atom picture of the interferometer (which approximates well the ‘enhanced’ scenario considered in this work where atomic cloud inhomogeneities remain small). The motion of the atom is described by Eq. 2 333Consistently with the rest of this paper, we still neglect here the Coriolis force due to the Earth’s rotation.. As the associated Lagrangian is quadratic in position and velocity, the interferometric phase shift can be computed using the mid-point-theorem Antoine and Bordé (2003), where (resp. ) is the laser wavevector along arm I (resp. II) of the interferometer and the average atomic position at pulse . The later can be obtained by solving Eq. 2. In the following, we will neglect variations of the fields , , and experienced by the atom during the interferometric sequence. Following the approach proposed in de Castanet et al. (2024), that we extend here to arbitrary rotations, an approximate polynomial solution to Eq. 2, truncated to third order, is sought for. Denoting the atom’s acceleration, its initial velocity and its initial position, all taken in the sensor frame, and rewriting for convenience, the phase shift is obtained as:
where is the momentum splitting of the interferometer and its interrogation time.
@@
where (resp. ) is the gravity field at the location of interferometer A (resp. B), (resp. ) is the acceleration of interferometer A (resp. B) in the Earth’s frame, and (resp. ) are the initial vertical position and velocity coordinates of the atom in interferometer A (resp. B) expressed in the sensor frame.
B.1 Requirements based on phase shift
@@
terms 3 to 6 in Eq. 6 can be evaluated as . We note here that the first term is always positive, meaning that constant rotations induce a positive bias on gravity gradient measurements, as observed for instance on Figs. 4-5. A sufficient condition for both these terms to remain bounded below is, for ms:
@@
Terms 7 to 9 in Eq. 6 involve, for each interferometer, the initial velocity of the atom in the interferometer. Assuming a typical thermal velocity, , the requirement that these terms remain smaller than would yield a stringent constraint on transverse temperature, e.g. nK if using and (which are 10 times lower than the upper bounds in Eq. 8) and 444The constraint on z-temperature is not stringent here, for rotations rates given by Eq. 8.. In situations where this condition breaks, such as in the ”moderate” scenario considered in this work, the interplay between platform rotation and atomic thermal velocities is therefore expected to contribute a phase shift. An exact calculation would however require a proper treatment of the thermal cloud of atoms, as done in this work, going beyond the simplified derivation presented in this Appendix.
B.2 Other considerations: effect of pulse infidelities
The above requirements are based solely on a simplified, single-atom, phase shift calculation.
@@ B.2 Other considerations: effect of pulse infidelities
An important one is the effect of platform accelerations. An acceleration of the platform does not cause any differential phase shift as it is common to both interferometers, but still causes, in each separate interferometer, pulse infidelity leading to a loss of fringe contrast and an apparent phase shift. This includes both longitudinal accelerations, which can induce additional Doppler shifts, and transverse accelerations, which cause the atoms to drift off-centred in the beam under the action of the centripetal force.
A simple, rough condition for this latter effect to remain negligible is that the atoms should remain within the beam during the full interrogation sequence, with the transverse acceleration of the considered interferometer, which typically gives for ms and a 1-mm beam radius. Importantly though here, would depend not only on the platform acceleration, but also on the platform rotation rate and the distance between the considered interferometer and the platform centre of rotation, through a relation similar to Eq. 7. Sensor tilts, which accentuate the atoms getting off-centered as they free-fall in the beam, and other inertial forces, which non-trivially affect atomic trajectories, can make this criterion significantly more involved - not to mention further complexity arising from atomic cloud inhomogeneities which lead to atom-dependent trajectories. An exact treatment requires as accurate a model as the one presented in this work, to account for the interplay between non-inertial atomic trajectories and atom-light interactions.
Going beyond the calculation presented in this Appendix, our atom interferometer model allows for an exact treatment: (i) not relying on an approximate solution for the atomic external dynamics; (ii) capturing couplings between the ballistic motion and the quantum dynamics of atoms at the level of each atom-light interaction - and not only at the phase shift level; (iii) capturing inter-atom dephasings and cross-couplings between the inhomogeneous cloud, inertial forces and light pulses; (iv) accounting for variations in inertial fields between ellipse shots, rather than relying on a phase shift obtained at fixed external fields.

### v2

@@ Appendix B Estimation of systematic effects and mitigation requirements
We consider here a simplified single-atom picture of the interferometer (which approximates well the ‘enhanced’ scenario considered in this work where atomic cloud inhomogeneities remain small). The motion of the atom is described by Eq. 2 222Consistently with the rest of this paper, we still neglect here the Coriolis force due to the Earth’s rotation.. As the associated Lagrangian is quadratic in position and velocity, the interferometric phase shift can be computed using the mid-point-theorem Antoine and Bordé (2003); Overstreet et al. (2021), where (resp. ) is the effective laser wavevector at pulse along arm I (resp. II) of the interferometer and the average atomic position at pulse . The latter can be obtained by solving Eq. 2, where and represent the acceleration and rotation vector of the sensor in the Earth’s frame. In the following, we assume that the sensor rotation vector varies at most linearly with time, and we neglect variations of the fields and experienced by the atom during the interferometric sequence. This notably means that for simplicity in this Appendix, gravity gradient effects within each individual interferometer, which are captured by our numerical model via the term in Eq. 2, are neglected in this calculation; a discussion of their effects is however provided in Sec. B.2. Following the approach proposed in d’Armagnac de Castanet et al. (2024), that we extend here to arbitrary rotations, we seek an approximate polynomial solution to Eq. 2 for , truncated to third order in time. We denote respectively , and the atom’s initial position, velocity and acceleration at the time of the first pulse; all vectors are defined in the sensor frame and expressed in a coordinate system attached with the sensor body where x is longitudinal, y is lateral and z is vertical. We assume the lasers to be strapdown with the sensor so that their wavevectors rotate along with it and the effective wavevector remains fixed in the sensor frame. Denoting with the vertical unit vector in the sensor’s frame, and rewriting for convenience, the phase shift is obtained as:
where is the interrogation time.
@@
where (resp. ) is the gravity field at the location of interferometer A (resp. B), (resp. ) is the acceleration of interferometer A (resp. B) in the Earth’s frame, and (resp. ) are the vertical position and velocity coordinates of the atom in interferometer A (resp. B) at the time of the first pulse, expressed in the sensor frame.
B.1 Requirements based on the differential phase shift Eq. 6
@@
where is the separation vector between the two interferometers A and B [ in sensor body axes], terms 3 to 5 in Eq. 6 can be evaluated as . We note here that the first term is always positive, meaning that constant rotations induce a positive bias on gravity gradient measurements, as observed for instance on Figs. 5-6. A sufficient condition for both these terms to remain bounded below is, for ms:
@@
Terms 6 to 8 in Eq. 6 involve, for each interferometer, the initial velocity of the atom at the time of the first pulse. Assuming a typical thermal velocity, , the requirement that these terms remain smaller than yields a stringent constraint on transverse temperature, which critically depends on the sensor rotation rate and rotational acceleration; as an example, for , assuming and (so that we are well into a regime where Eq. 8 is fulfilled and purely rotational effects are negligible), we would find nK 333The constraint on z-temperature is not stringent here, for rotations rates given by Eq. 8.. In situations where this condition breaks, such as in the ”moderate” scenario considered in this work, the interplay between platform rotation and atomic thermal velocities is therefore expected to contribute a phase shift. An exact calculation would however require a proper treatment of the thermal cloud of atoms, as done in this work, going beyond the simplified derivation presented in this Appendix.
B.2 Further considerations
Effects of gravity gradients:
The above calculation has neglected variations of the gravity field inside each individual interferometer. At lowest order in the gravity gradient, the phase shift of an interferometer in the presence of a gravity gradient is given by Peters et al. (2001). When considering a differential measurement between two interferometers A and B, the first two terms produce the gradiometric signal of interest , while the two others give rise, at lowest order in the gravity gradient, to corrections to the differential phase shift
Assuming thermal velocities for and , and for ms, we find that the first term will remain smaller than provided the z-temperature is below .
In turn, the requirement that the second term is smaller than sets a constraint on the sensor rotation rate (by virtue of Eq. 7), , which is however much less stringent than Eq. 8.
An more exact treatment of gravity gradient effects capturing higher couplings with sensor motion and tilts would require solving Eq. 2 in the presence of a inhomogeneous gravity field. While this goes beyond the scope of this Appendix, this is captured in our numerical interferometer model.
Effects of pulse infidelities:
All the above requirements are based solely on a simplified, single-atom, phase shift calculation.
@@ B.2 Further considerations
An important one is the effect of platform accelerations. An acceleration of the platform does not cause any differential phase shift as it is common to both interferometers, but still causes, in each separate interferometer, pulse infidelity leading to a loss of fringe contrast and an apparent phase shift. This includes both longitudinal accelerations, which induce additional Doppler shifts, and transverse accelerations, which cause the atoms to drift off-centred in the beam under the action of the centripetal force.
As a rough estimate, longitudinal accelerations will typically start to play a role when the Doppler shift at the final pulse, , is of the order of the pulse bandwidth , which gives for the ”moderate” sensor specifications and for the ”enhanced” sensor specifications.
As for transverse accelerations, a simple condition is that the atoms should remain within the beam during the full interrogation sequence, with the transverse acceleration of the considered interferometer, which typically gives for ms and a 1-mm beam radius. Importantly though here, would depend not only on the platform acceleration, but also on the platform rotation rate and the distance between the considered interferometer and the platform centre of rotation, through a relation similar to Eq. 7. Sensor tilts, which accentuate the atoms getting off-centered as they free-fall in the beam, and other inertial forces, which non-trivially affect atomic trajectories, can make these criteria significantly more involved - not to mention further complexity arising from atomic cloud inhomogeneities which lead to atom-dependent trajectories. An exact treatment requires as accurate a model as the one presented in this work, to account for the interplay between non-inertial atomic trajectories and atom-light interactions.
In summary, going beyond the calculation presented in this Appendix, our atom interferometer model allows for an exact treatment: (i) not relying on an approximate solution for the atomic external dynamics; (ii) accounting for variations in the experienced external fields, e.g. , during the interferometric sequence; (iii) capturing couplings between the ballistic motion and the internal dynamics of atoms at the level of each atom-light interaction - and not only at the phase shift level; (iv) capturing inter-atom dephasings and cross-couplings between the inhomogeneous cloud, inertial forces and light pulses; (v) accounting for variations in inertial fields between ellipse shots, rather than relying on a phase shift obtained at fixed external fields.